# Three-Wave Interaction Grating Coupler with Sub-Decibel Insertion Loss at Normal Incidence

### v1

@@ 1 Introduction
Silicon photonics is making inroads in optical systems ranging from optical communication [1, 2], spectroscopy [3, 4, 5], and quantum optics [6, 7] to astro-photonics [8, 9, 10]. The gold standard for coupling into silicon photonic single mode waveguides from free space or standard single-mode optical fiber is edge coupling, which can achieve both high coupling efficiency and large bandwidths [11, 12]. However, edge coupling has several drawbacks. It requires additional preparation of the facets of the silicon photonics waveguides and poses challenges to wafer-level testing, requiring specialized equipment [13]. Additionally, in an edge coupling configuration the mode field diameter (MFD) of the integrated waveguide is limited by the thickness of the lower cladding which is typically 2 m to 3 m thick. As a result, efficient mode matching to edge couplers requires lens tapered fibers which are expensive, fragile, and impose strict lateral alignment tolerances when compared to larger MFD single mode fibers [14, 11, 12].
@@ 2 Modeling
Together equations 3, 4, and 5 completely parameterize the geometry of a TWIG coupler of a given material stack with selected etch depths , , and . The material stack available to us for this work uses a 725 m thick silicon handle wafer with a 2 m thick thermally grown buried oxide layer. The silicon waveguide layer is a standard 220 nm thick device layer. An additional 2.5 m of oxide cladding is deposited above the waveguide layer following patterning. For the etch depths we have selected = = 220 nm. We have set these two depths equal such that they can be realized in the same process step and minimize the number of lithography steps needed. We have decided to fully etch through the waveguiding layer to remove any possibility of an over-etching error. The third scattering site is assigned a partial etch depth of = 70 nm. It must be noted that these etch depths were assigned somewhat arbitrarily and that the TWIG geometry can accommodate any number combinations for etch depths.
@@ 3 Optimization
To ensure a reasonable balance between bandwidth and coupling efficiency, the figure of merit is defined as the average coupling efficiency over several wavelengths of interest. Prior work has shown that the method is effective at improving bandwidth with small impact to the peak coupling efficiency [36]. Additionally, we impose two design restrictions with regards to the minimum feature size of the TWIG coupler based on the tolerances provided by Applied Nanotools. It is enforced that all features with a full etch depth of 220 nm have a minimum critical dimension (etch width) of 70 nm and that all features with a partial etch depth of 70 nm have a minimum critical dimension of 160 nm. These restrictions are enforced during the optimization via a penalty term appended to the FOM.
@@
From this optimized design we generate two variations of the TWIG coupler: a one-dimensional grating that only varies along the direction of propagation and a two-dimensional focusing grating coupler. The one-dimensional grating coupler maintains a constant width of 13 m over the length of the grating and employs a 150 m long adiabatic taper to transition between the width of the grating and a single mode waveguide. The two-dimensional grating coupler varies radially, introducing a curved phase profile which focuses the coupled light to a single mode waveguide at the origin of the radial etches [15]. The radial etches span an angle of and begin after an initial offset of 24.35 m in order to achieve an approximate 10.4 m MFD. The two-dimensional focusing coupler reduces the length required to transition from the grating mode to a single mode waveguide. The required footprint of the focusing TWIG is 41 m 16 m while the one-dimensional grating requires a footprint of 150 m 13 m. Figure 2d and 2e displays the two geometries with the simulated grating modes overlaid.
@@ 5 Experimental Results
The devices described above were submitted to the commercial foundry Applied Nanotools for fabrication. Applied Nanotools employs electron beam lithography to reliably achieve feature sizes down to 70 nm, ideal for low volume applications. Similar and smaller feature sizes can be achieved using optical lithography techniques for larger volumes as well. Figure 4 displays SEM micrographs of the fabricated couplers.
Our test structures consist of an input TWIG coupler, a straight single mode waveguide, 500 nm 200 nm, with a length of 1.03 cm, and an identical TWIG coupler rotated to act as an output coupler. We use an HP81680A tunable laser which operates in and around the c-band as our input source. The laser is fiber coupled to the chip using polarization maintaining single mode fiber. The polarization maintaining fiber has been cleaved and is held within a brass ferrule, which is itself held within a rotatable mount. The entire ferrule may be rotated to adjust the input polarization relative to the TWIG coupler. A cleaved SMF28 fiber is used at the output coupler before detection at an HP81531A photodiode. Alignment of the input and output fibers is managed by separate Thorlabs Nanomax 600 series 6 axis stages. Additionally, an index matching fluid (n = 1.444) has been used here to alleviate Fresnel reflections between the cladding/air and fiber/air interfaces.
@@
To properly extract the efficiency of each TWIG coupler from an optical transmission measurement, we must first account for the propagation losses associated with the single mode waveguide. The experimental coupling efficiency is given by the function below where is the detected power, is the input power, and is the transmission efficiency of the waveguide.
We have estimated the propagation losses of our single mode waveguides from the Fabry-Perot resonances observed during the measurement of the highest performing one-dimensional TWIG coupler [38]. The Fabry-Perot resonances observed are a result of weak back-reflections waveguide/coupler interface. The transmission spectra is given in figure 5 where a maximum transmission of 47.8 is observed. The free spectral range of the Fabry-Perot resonance is 0.029 nm, which corresponds to a cavity length of 1.03 cm, the separation between the two gratings. To estimate the propagation losses, we perform a fitting of the measured Fabry-Perot spectra to the analytical solution of a Fabry-Perot cavity with propagation losses. To constrain the parameter space for this fitting, the power reflectivity of the gratings is assumed to be 4 based on FDTD simulations of the optimal device. Using this method, we estimate a propagation loss of -1.35 dB/cm which is in good agreement with the average value reported by Applied Nanotools.
@@
The one-dimensional TWIG coupler plotted in figure 6 has been designed with -5 nm etch width bias such that it compensates for a 5 nm over etch. Additionally, this device has been designed for a -10 nm alignment error correction. The highest performing focusing TWIG coupler has been designed with a -10 nm etch width bias correction and a -10 nm alignment error. The difference in etch width bias between these two highest performing devices may be a result of local etch rate variations caused by the difference in geometry.
Figure 7 shows the measured coupling spectra of several of the remaining focusing TWIG couplers, displaying the impact of each type of fabrication variation. The wavelength resolution of the sweep here has been reduced for the convenience of viewing the closely overlapping measurements. We see good agreement between the experimental data and simulation with regard to both types of fabrication errors. We can see from panel 7a that a positive etch width bias causes the spectra to broaden and shift toward shorter wavelength, while reducing peak coupling efficiency. We observe that -10 nm correction has the best performance among the devices with no overlap error. This indicates that the true etch width bias is close to -10 nm. From panel 7b we can see that regardless of alignment error each of the devices performs similarly demonstrating robustness to this type of fabrication variation.

### v2

@@ 1 Introduction
Silicon photonics is making inroads in optical systems ranging from optical communication [1, 2], spectroscopy [3, 4, 5], and quantum optics [6, 7] to astro-photonics [8, 9, 10]. The gold standard for coupling into silicon photonic single mode waveguides from free space or standard single-mode optical fiber is edge coupling, which can achieve both high coupling efficiency and large bandwidths [11, 12]. However, edge coupling has several drawbacks. It requires additional preparation of the facets of the silicon photonics waveguides and poses challenges to wafer-level testing, requiring specialized equipment [13]. Additionally, in an edge coupling configuration the mode field diameter (MFD) of the integrated waveguide is limited by the thickness of the lower cladding which is typically 2 m to 3 m thick. As a result, efficient mode matching to edge couplers often requires either complex processing to allow for suspended silicon edge couplers [14] or lens tapered fibers, which are expensive, fragile, and impose strict lateral alignment tolerances when compared to larger MFD single mode fibers [15, 11, 12].
@@ 2 Modeling
Together equations 3, 4, and 5 completely parameterize the geometry of a TWIG coupler of a given material stack with selected etch depths , , and . The material stack available to us for this work uses a 725 m thick silicon handle wafer with a 2 m thick thermally grown buried oxide layer. The silicon waveguide layer is a standard 220 nm thick device layer. An additional 2.5 m of oxide cladding is deposited above the waveguide layer following patterning. For the etch depths we have selected = = 220 nm. We have set these two depths equal such that they can be realized in the same process step and minimize the number of lithography steps needed. We have decided to fully etch through the waveguiding layer to remove any possibility of an over-etching error. The third scattering site is assigned a partial etch depth of = 70 nm. It must be noted that these etch depths were assigned somewhat arbitrarily and that the TWIG geometry can accommodate any number combinations for etch depths [37].
@@ 3 Optimization
To ensure a reasonable balance between bandwidth and coupling efficiency, the figure of merit is defined as the average coupling efficiency over several wavelengths of interest. We have chosen to discretize the FOM over only a few wavelengths of interest to reduce the computational load of each iteration in the optimization. Prior work has shown that the method is effective at improving bandwidth with small impact to the peak coupling efficiency [37]. Additionally, we impose two design restrictions with regards to the minimum feature size of the TWIG coupler based on the tolerances provided by Applied Nanotools. It is enforced that all features with a full etch depth of 220 nm have a minimum critical dimension (etch width) of 70 nm and that all features with a partial etch depth of 70 nm have a minimum critical dimension of 160 nm. These restrictions are enforced during the optimization via a penalty term appended to the FOM.
@@
It can be seen that several of the local periods of the optimized design share similarities with designs discussed in [18], which investigates blazed gratings employing a nanopillar and L-shaped structure. It is the case that such a geometry is a subset of the TWIG coupler which requires that the partially etched sections overlap with the fully etched sections. Here, we solve and optimize a more generalized geometry, allowing for more degrees of freedom with which to satisfy the phase matching conditions for high directionality.
From this optimized design, we generate two variations of the TWIG coupler: a one-dimensional grating that only varies along the direction of propagation and a two-dimensional focusing grating coupler. The one-dimensional grating coupler maintains a constant width of 13 m over the length of the grating and employs a 150 m long adiabatic taper to transition between the width of the grating and a single mode waveguide. The two-dimensional grating coupler varies radially, introducing a curved phase profile which focuses the coupled light to a single mode waveguide at the origin of the radial etches [16]. The radial etches span an angle of and begin after an initial offset of 24.35 m in order to achieve an approximate 10.4 m MFD. The two-dimensional focusing coupler reduces the length required to transition from the grating mode to a single mode waveguide. The required footprint of the focusing TWIG is 41 m 16 m while the one-dimensional grating requires a footprint of 150 m 13 m. Figure 2d and 2e displays the two geometries with the simulated grating modes overlaid.
@@ 5 Experimental Results
The devices described above were submitted to the commercial foundry Applied Nanotools for fabrication. Applied Nanotools employs electron beam lithography to reliably achieve feature sizes down to 70 nm, ideal for low volume applications. As these designs require feature sizes below 100 nm, they are not compatible with the 193 nm process node typically used for integrated photonics multi-project wafer runs. However, increasingly advanced optical lithography is being pursued for integrated photonic applications, reducing the minimum feature sizes provided by commercial foundries. Figure 4 displays SEM micrographs of the fabricated couplers.
Our test structures consist of an input TWIG coupler, a straight single mode waveguide, 500 nm 220 nm, with a length of 1.03 cm, and an identical TWIG coupler rotated to act as an output coupler. We use an HP81680A tunable laser which operates in and around the c-band as our input source. The laser is fiber coupled to a polarization maintaining circulator, which enables us to monitor back reflections from the structure. The output of the circulator is coupled to the chip using a polarization maintaining fiber that has been cleaved and is held within a brass ferrule. The brass ferrule itself is held within a rotatable mount enabling us to tune the input polarization relative to the TWIG coupler. A cleaved SMF28 fiber is used at the output coupler before detection at an HP81531A photodiode. Alignment of the input and output fibers is managed by separate Thorlabs Nanomax 600 series 6 axis stages. Additionally, an index matching fluid (n = 1.444) has been used here to alleviate Fresnel reflections between the cladding/air and fiber/air interfaces.
@@
To extract the efficiency of each TWIG coupler from an optical transmission measurement, we must first account for the propagation losses associated with the single mode waveguide. The experimental coupling efficiency is given by the function below where is the detected power, is the input power, and is the transmission efficiency of the waveguide.
We have estimated the propagation losses of our single mode waveguides from the Fabry-Perot resonances observed during the measurement of the highest performing one-dimensional TWIG coupler [39, 40]. The Fabry-Perot resonances observed are a result of weak back-reflections at the waveguide/coupler interface. The measured transmission spectrum is given in figure 5(a) where a maximum transmission of 47.8 is observed at a wavelength of 1555.5 nm. The free spectral range of the Fabry-Perot resonance is 0.029 nm, which corresponds to a cavity length of 1.03 cm, the separation between the two gratings.
The measured reflection spectrum is given in figure 5(b) exhibiting a complimentary Fabry-Perot resonance with a similar free spectral range. At short wavelengths, we observe minimal back reflections on the order of -20 dB which increase sharply as the wavelength increases beyond 1560 nm. The measured back reflections around the peak coupling wavelength are observed to fluctuate between 0.1 and 4.5. This is of similar order of magnitude to back reflections measured for other techniques used to reduce reflections in perfectly vertical grating couplers [41]. In this work back reflections were not considered in the FOM used for optimization. In future work, the FOM may be adjusted to explicitly minimize back reflections albeit with some trade off to the peak coupling efficiency.
To estimate the propagation losses, we perform a fitting of the measured transmission spectra to the analytical solution of a Fabry-Perot cavity with propagation losses. Within this fitting, the power reflectivity of the gratings is left as a fitting parameter. The optimal fitting parameters suggest a propagation loss of -1.35 dB/cm and a power reflectivity of 4.35, which are in good agreement with the average value reported by Applied Nanotools (-1.2 dB/cm) and the measured back reflections.
@@
Both devices exhibit an abrupt decrease in performance around 1528 nm. We believe this to be a measurement artifact associated with our testing station. This is indicated by the observation that each structure exhibits a similar drop at the same wavelength, 1528 nm, regardless of the shift to the peak coupling wavelength.
The one-dimensional TWIG coupler plotted in figure 6 has been designed with a -5 nm etch width bias such that it compensates for a 5 nm over-etch. Additionally, this device has been designed for a -10 nm alignment error correction. The highest performing focusing TWIG coupler has been designed with a -10 nm etch width bias correction and a -10 nm alignment error. The difference in etch width bias between these two highest performing devices may be a result of local etch rate variations caused by the difference in geometry.
Figure 7 shows the measured coupling spectra of several of the remaining focusing TWIG couplers, displaying the impact of each type of fabrication variation. The wavelength resolution of the sweep here has been reduced for the convenience of viewing the closely overlapping measurements. We see good agreement between the experimental data and simulation with regard to both types of fabrication errors. We can see from panel 7a that a positive etch width bias causes the spectra to broaden and shift toward shorter wavelengths while reducing peak coupling efficiency. We observe that a -10 nm correction exhibits the best performance among the devices with no overlap error. This indicates that the true etch width bias is close to -10 nm. From panel 7b we can see that, regardless of alignment error, each of the devices performs similarly, demonstrating robustness to this type of fabrication variation.
Figure 8a provides the measured back reflection spectra that result from an etch width bias in the absence of alignment errors. We observe an identical trend to that of the coupling efficiency. A positive etch width bias results in an increase in the measured back reflection intensity as well as a blue-shift of the center wavelength. Figure 8b displays the measured back reflection that results from alignment errors in the absence of etch width bias corrections. We similarly observe that the measured back reflection varies little for alignment errors on the order of 10 nm.
For potential etch width errors spanning 10 nm, we observe that the peak wavelength shifts by approximately 30 nm and the peak coupling efficiency is reduced from 85 to 50. For the same range of errors, the measured back reflections undergo a similar shift. For a +10 nm etch width bias, we observe back reflections as high as -20 (-7 dB) for wavelengths longer than 1540 nm. The observed phenomenon is likely caused by the gradual violation of the phase matching conditions required for high directionality [37]. These results suggest fairly stringent tolerances with regard to etch width bias; however, we observe that the device is robust to variations in the alignment between the full etch and partial etch steps. In future work, additional terms may be appended to the FOM to reduce the variability of device performance with respect to etch width bias.